# Bounds and Bugs: The Limits of Symmetry Metrics to Detect Partisan Gerrymandering

### v1

@@ 2 Theoretical Results
Note from Theorems 1 and 2, that when $S\geq\frac{1}{2}$, both the Mean-Median Difference and Partisan Bias cannot be zero for $V<\frac{1}{2}$, and that both the Mean-Median Difference and Partisan Bias can be zero for any $V$ with $\frac{1}{2}\leq V<\frac{3}{4}$. We can deduce the remainder of Corollary 1 by setting the lower bound equal to 0 in the region $\frac{3}{4}\leq V\leq 1$.
@@
Note that, by symmetry, we have all $(V,S)$ pairs for which there exists constructed election data with vote share $V$, seat share $S$, and Partisan Bias 0. This result shows that, when turnout in each district is equal, a very wide array of $(V,S)$ pairs can result in $MM=0$ and $PB=0$. The region consisting of all $(V,S)$ for which there is constructed election data with $MM=0$ and $PB=0$ can be seen in Figure 6.
@@
This is different from the set of $(V,S)$ for which there is constructed election data with turnout equal in all districts and $\text{Declination}=0$ [7] (see Figure 7), and also quite different from the set of $(V,S)$ for which there is constructed election data with turnout equal in all districts and $EG=0$ [34] (see Figure 8). From these Figures, we can deduce that the Partisan Bias and Mean-Median Difference are the least restrictive measures in terms of which $(V,S)$ pairs are allowed to potentially have a metric value which indicates that no gerrymandering has occurred.
@@
Note that Corollary 1 assumes that turnout in each district is equal, which is never the case in real elections. This raises the question: if turnout is unequal, say the maximum turnout divided by minimum turnout is $C$, what $(V,S)$ pairs allow for Partisan Bias and Mean-Median Difference 0? The following result answers this question.
@@
Note that Theorem 3 gives the lower bound of $\frac{1}{2}$ when $C=1$ (when turnout in all districts is equal), which is what we already know to be true from Corollary 1. For the upper bound, when $C=1$ we again get the result of Corollary 1. Specifically, setting $C=1$, Theorem 3 gives $V^{*}\leq\frac{2-S}{3-2S}$, which is the result of solving the equation $S=\frac{3V^{*}-2}{2V^{*}-1}$ (from Corollary 1) for $V^{*}$. When turnout is unequal, Theorem 3 expands the pairs $(V,S)$ for which there is election data with vote share $V$, seat share $S$, and Partisan Bias/Mean-Median Difference values of 0. For example, Figure 9 shows how this region expands for different values of $C$. As $C\to\infty$, the region expands to fill the $[0,1]\times[0,1]$ square.
@@
Theorem 3 is also proven in Section A.
3 Empirical Results: Difference between the Mean-Median and Partisan Bias
From Section 2, we see that the set of $(V,S)$ pairs for which there is election data giving $MM=0$ is the same as the set of $(V,S)$ pairs for which there is election data giving $PB=0$. However, the actual values of the Mean-Median Difference and Partisan Bias can be different on a single election. In this section, we empirically study the $(MM,PB)$ pairs found in many US maps.
All of the data for this study were collected from Dave’s Redistricting App [12] between April 19 and April 21, 2024. For each state, all districting maps available were collected; this typically consisted of Congressional districts, State House districts, and State Senate districts [^4] Several state house and state senate districting maps have districts with extremely different populations, for various reasons. For example, the Vermont state senate has each district being a multimember district, but with a range of 1-3 members per district. Those maps were simply included as-is, and did not give the most extreme values of $MM-PB$ that we observed., with exceptions for states that have one congressional district (consisting of the whole state) and/or a single state legislative map. (If a state has a single state legislative map, that map was included in the “State Senate” maps). The maps collected were those listed as “current maps” according to DRA, and using the partisan data on the site (which is a composite of several election results).
We can see a scatter plot of $(MM,PB)$ pairs for all of the maps analyzed in Figure 10; there are 44 Congressional maps, 45 State House maps, and 50 State Senate maps for a total of 139 maps all together. Positive values are intended to indicate a map that benefits the Democratic party; negative values are intended to indicate a map that benefits the Republican party.
From these scatterplots, we can see that the Mean-Median Difference and Partisan Bias values can be quite different in real-world data. Recall from Section 1.5 that the supremal theoretical difference between the Mean-Median Difference and Partisan Bias is 0.5. The largest difference seen in our data was 0.33, which was achieved on South Carolina’s congressional map. Other maps with a large difference (in absolute value) were Massachusetts congressional (0.25), Indiana Congressional (0.23), Wisconsin Congressional (0.20), Iowa Congressional (-0.22), and Nevada Congressional (-0.22). The only states having both Mean-Median Difference and Partisan Bias values of 0 were those where it is guaranteed to have $MM=PB=0$: states with two congressional districts (Hawaii, Idaho, Maine, Montana, New Hampshire, Rhode Island, and West Virginia).
Recall that values far from 0 for these metrics is intended to indicate a greater degree of gerrymandering, while the sign is intended to indicate which party was gerrymandered. As stated in Section 1.5, the sign of the two metrics cannot differ, although the values can. These results indicate that the Mean-Median Difference and Partisan Bias metrics can disagree quite drastically as to the degree of gerrymandering happening within a state.
We also note that the sign of these metrics (positive indicating a map benefitting Democrats and negative indicating a map benefitting Republicans) on certain states may be surprising to some readers. For example, the Partisan Bias and Mean-Median Difference suggest that the Massachusetts congressional map is potentially a Republican gerrymander. The seats-votes curve for that map was in Figure 1; we encourage the reader to explore these plots and consider how frequently the sign of these partisan symmetry metrics are incorrect (the incorrectness of sign was also addressed in [10]).
Figure 10: $(MM,PB)$ pairs for all for State House, State Senate, and Congressional districts, all states. Data gathered from Dave’s Redistricting App between April 19 and April 21, 2024, for maps listed as “current maps” at the time. Top image allows points to be more easily distinguished, bottom image has the same scale on the vertical and horizontal axes. Dotted lines correspond to $MM=PB$.
4 Empirical Results: Accuracy in a Neutral Ensemble and Reliability
From Section 2, we see that both the Mean-Median Difference and Partisan Bias allow for a wide variety of seats/votes pairs corresponding to election data with metric value of 0. This suggests that potentially these metrics are less powerful at detecting gerrymanders. Here we provide empirical results that support this suggestion.
4.1 Outlier Analysis
One of the techniques of detecting gerrymandering that is most widely accepted by the mathematical and computational community is outlier analysis. Through this technique, a neutral ensemble of typically tens of thousands of redistricting maps is created, to which the proposed map is compared. If the proposed map looks like an outlier compared to the other maps in the ensemble, this could suggest the presence of gerrymandering. Several research groups implement different varieties of this outlier analysis; we refer the reader to the Metric Geometry Gerrymandering group [20] or the Quantifying Gerrymandering group [21] for examples.
This outlier analysis can use any kind of metric, from the Mean-Median Difference or Partisan Bias, GEO metric, Efficiency Gap, to the simple metric of number of districts won by a particular party. It is a natural question to ask: does the Mean-Median Difference or Partisan Bias seem to do a good job when it comes to neutral ensembles? That is, does it agree with the other metrics in determining which maps are gerrymanders? Or does it have positive errors (detecting a gerrymander where there is none) or negative errors (missing a clear gerrymander)? Note that we consider the number of districts won to be the foundation of a partisan gerrymander (see the discussion in Section 1.2).
This question was answered for the Mean-Median Difference in [30], which showed that the Mean-Median Difference commits both positive and negative errors. Here we show the same for the Partisan Bias.
For the purposes of this study, we used the exact same 18 maps that were studied by the authors of [30], but use three elections for each state ([30] considered one election per state). Those 18 maps corresponded to 3 maps per state in 6 states which have different political geographies: Massachusetts, Michigan, Oklahoma, Oregon, Pennsylvania, and Texas. Data for each state was taken from the Metric Geometry Gerrymandering Group’s GitHub repository containing data from the maps created after the 2010 Census cycle [19]. For each state, we used the most recent Gubenatorial, Presidential, and US Senate election that was available in the datset at [19], with the exception of Oklahoma, which had no Presidential or US Senate election data (we chose the 2018 Attorney General and 2018 Lieutenant Governor elections instead). We also use the same Python library of gerrychain [28] that was used in [30]; gerrychian implements a Markov Chain Monte Carlo process in order to create the neutral ensemble of potential redistricting maps. We used a population deviation of 5% for congressional maps and 11% for state house/state senate maps in all of our ensembles, and we created an ensemble of size 20,000 (also mimicking what was used in [30]).
A sample result can be seen in Figure 11. Please note that the GEO metric is somewhat different from other metrics in that large values of a Democratic GEO score indicate a map that is less beneficial to the Democratic party, whereas for all the other metrics a large value indicates a map that is more beneficial to the Democratic party. For this reason, we display the negated Democratic GEO score in our histograms, so that for all histograms smaller values are less beneficial to the Democratic party. We also note that the creators of the GEO score emphasized that both the Democratic and Republican GEO scores should be used together (not separately) to evaluate a map [6].
Figure 11: Neutral ensembles for Massachusett’s State House districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric.
Here we can see that all of the metrics with all elections considered suggest the same thing: the enacted map looks fairly typical compared to the maps in the neutral ensemble. However, three of the 18 considered maps revealed that the Partisan Bias, like the Mean-Median Difference, commits both positive and negative errors. Consider Figures 12 and 13.
Figure 12: Neutral ensembles for Massachusett’s Congressional districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric. Declination not included because it is not defined for nearly all maps. Partisan Bias suggests that this map is a gerrymander on all election data, unlike all other metrics.
Figure 13: Neutral ensembles for Texas’s Congressional districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric. Partisan Bias suggests that this map is a gerrymander on all election data, unlike all other metrics. Both GEO scores paired together raise a possibility of gerrymandering on the GOV 14 data.
In Figures 12 and 13 we can see that both the Massachusetts and Texas Congressional districting maps are flagged as outliers by the Partisan Bias metric on all election data, unlike any other metric. Additionally, research has clearly shown that, while MA has all 9 districts going to the Democratic party, it is essentially impossible to draw a map in which Republicans can win a single district [11] (indicating that the Massachusetts Congressional map is not a gerrymander).
Alternatively, we can see in Figure 14 that the Partisan Bias does not flag Pennsylvania’s 2011 congressional map as an outlier on the two election data that all other metrics do (aside from the Mean-Median Difference, which does not detect the gerrymander on the T16SEN data). This is despite the fact that the other metrics do flag it as an outlier (except for the Mean-Median Difference!) and the wide acceptance that that map was indeed a gerrymander (including by the Pennsylvania Supreme Court [25]).
Figure 14: Neutral ensembles for Pennsylvania’s Congressional districting plan. The red line corresponds to the enacted plan’s value for the corresponding metric. All metrics suggest gerrymandering on T16SEN and GOV14 data, but not on T16PRES data, except for Partisan Bias (which never detects gerrymandering) and Mean-Median Difference (which does not detect it on the T16SEN data). Recall that both GEO scores must be used together.
4.2 Short Burst Analysis
The authors of [30] also used a Short Burst analysis to test the “gameability” of some metrics intended to detect gerrymandering (specifically, the GEO Metric, Mean-Median Difference, Declination, Efficiency Gap). They define a metric as “gameable” if the metric value can stay within acceptable bounds [^5] For a metric $m$, they considered “acceptable bounds” to be $-0.16\inf(m)\leq m\leq 0.15\sup(m)$, based on the bounds recommended by the creators of the Efficiency Gap[31]., while allowing a map with the most extreme number of Democratic-won (or Republican-won) districts that can be achieved by that map.
@@ 4.2 Short Burst Analysis
In order to find such extreme maps, they use the Short Burst method introduced by Cannon et all in [8] [^6] The idea behind short bursts was suggested by Z. Schutzman at the Voting Rights Data Institute in the summer of 2018. In their article, Cannon et al used Short Bursts to find the largest number of majority-minority districts that a state can have in a single map. The short burst method takes a small number of steps in a Markov Chain (say 10 steps). At the end of that “short burst”, the map with the highest number of majority-minority districts is then the seed for the next short burst of size 10. This process is repeated thousands of times in an effort to find the maximum number of majority-minority districts achievable. In [8] Cannon et al showed that short bursts work better than biased walks in Markov Chains in order to find that maximum. Note that the Short Burst method can be thought of as a non-deterministic version of beam search on the metagraph [^7] The “metagraph” here is the graph whose nodes are maps. Two maps are adjacent if one can be obtained by the other through a single step in the Markov Chain constructing the ensemble of maps. In this case, that is through a single ReCom move in GerryChain. of redistricting maps, using depth-first search instead of breadth-first, and in which a random selection of maps near to the current map are explored.
In [30], Ratliff et al used this same technique, but in order to maximize the number of Democratic-won or Republican-won districts. We use the same technique here, on the same 18 maps that were discussed in Section 4.1 and using the same election data used in [30] [^8] We used a single election outcome (the same used in [30]) because the goal here is not to assess whether a map is an outlier, but just to assess bounds on the metric for maps with a fixed number of districts won.. Perhaps the most startling results in [30] were that, while the GEO Metric, Declination, and Efficiency Gap values tended to go up when the number of Democratic-won districts increased (indicating a map that benefits the Democratic party), the Mean-Median Difference values did not. The authors of [30] did not evaluate the Partisan Bias (likely because that metric is less frequently discussed in recent academic research), but we do so here. We show that the Partisan Bias shares the same uninformative behavior of the Mean-Median Difference.
For example, consider Figures 15 and 16.

### v3

@@ 4.4 The Special Case of Metric Value 0
From Theorems 1 and 3, we deduce that when $S\geq\frac{1}{2}$, both the Mean-Median Difference and Partisan Bias cannot be zero for $V<\frac{1}{2}$, and that both the Mean-Median Difference and Partisan Bias can be zero for any $V$ with $\frac{1}{2}\leq V<\frac{3}{4}$. We can deduce the remainder of Corollary 1 by setting the lower bound equal to 0 in the region $\frac{3}{4}\leq V\leq 1$.
@@
By symmetry, we have all $(V,S)$ pairs for which there exists constructed election data with vote share $V$, seat share $S$, and Partisan Bias 0. This result shows that, when turnout in each district is equal, a very wide array of $(V,S)$ pairs can result in $MM=0$ and $PB=0$. The region consisting of all $(V,S)$ for which there is constructed election data with $MM=0$ and $PB=0$ can be seen in Figure 5.
@@
This is different from the set of $(V,S)$ for which there is constructed election data with turnout equal in all districts and $\text{Declination}=0$ [9] (see Figure 6), and also quite different from the set of $(V,S)$ for which there is constructed election data with turnout equal in all districts and $EG=0$ [48] (see Figure 7). From these Figures, we can deduce that the Partisan Bias and Mean-Median Difference are the least restrictive metrics in terms of which $(V,S)$ pairs are allowed to potentially have a metric value which indicates that no gerrymandering has occurred.
@@
Corollary 1 assumes that turnout in each district is equal, which is never the case in real elections. This raises the question: if turnout is unequal, say the maximum turnout divided by minimum turnout is $C$, what $(V,S)$ pairs allow for Partisan Bias and Mean-Median Difference 0? The following result answers this question.
@@
Theorem 5 gives the lower bound of $\frac{1}{2}$ when $C=1$ (when turnout in all districts is equal), which is what we already know to be true from Corollary 1. For the upper bound, when $C=1$ we again get the result of Corollary 1. Specifically, setting $C=1$, Theorem 5 gives $V^{*}\leq\frac{2-S}{3-2S}$, which is the result of solving the equation $S=\frac{3V^{*}-2}{2V^{*}-1}$ (from Corollary 1) for $V^{*}$. When turnout is unequal, Theorem 5 expands the pairs $(V,S)$ for which there is election data with vote share $V$, seat share $S$, and Partisan Bias/Mean-Median Difference values of 0. For example, Figure 8 shows how this region expands for different values of $C$. As $C\to\infty$, the region expands to fill the entire $[0,1]\times[0,1]$ square! That is, for any $(V,S)$ with $10<V<$ 1 and $10<S<$ 1, there exists constructed election data131313with potentially extremely varied turnout between districts with $MM=0$, and the same can be said for $PB$.
The authors of [48] and [9] calculated similar regions for the Efficiency Gap and the Declination. For both of those metrics, the set of $(V,S)$ pairs for which the Efficiency Gap (Declination) can be 0 does not fill the entire $[0,1]\times[0,1]$ square, even when there are no restrictions on turnout. Again, this indicates that the Partisan Bias and Mean-Median Difference are the least restrictive metricss in terms of which $(V,S)$ pairs are allowed to potentially have a metric value which indicates that no gerrymandering has occurred, including when there are no restrictions on turnout.
@@
Theorem 5 is also proven in Appendix A.
5 Bugs
In Section 4, we stated Theorems (proved in Appendix A) that gave bounds on the values that the Mean-Median Difference and Partisan Bias could take on for each possible vote-share, seat-share pair $(V,S)$. We showed how one could use those results to see that the MM and PB do not allow for more extreme metric values when a more extreme number of districts is won. We also constructed election examples in which the metric values run counter to how these metrics are expected (and advertised) to act. Of course, the space of constructible examples may be quite different from the space of implemented redistricting maps. Here we give clear evidence that, on real maps with real data, the Mean-Median Difference and Partisan Bias cannot detect maps with an extreme number of districts won.
In order to test whether the Mean-Median Difference and the Partisan Bias have extreme values on maps with an extreme number of districts won, we need to produce maps with an extreme number of districts won. We do this with a recently developed hill-climbing method called the “Short Burst Method,” developed by Cannon et al [10]. This hill-climbing process uses GerryChain, which is a Python library that implements a well-researched and widely accepted Markov Chain Monte Carlo (MCMC) process to create an ensemble of potential redistricting maps [18, 40]. In [10], Cannon et al showed that this hill-climbing method worked better than a biased random walk [^14] A biased random walk is the most common method of using a Markov Chain to find local or global maxima; in this case, we are searching for redistricting maps with a near-maximum number of districts won by some party. at finding maps with an extreme number of majority-minority districts. Our application is a direct parallel; while Cannon et al were looking for maps with an extreme number of districts which were majority-minority, here we are looking for maps with an extreme number of districts which are majority-Democrat or majority-Republican.
5.1 Methods
@@
As stated above, we use the Short Burst Method to construct maps with an extreme number of districts won by a single party. The Short Burst Method takes a small number of steps in a Markov Chain (say 10 steps). At the end of that “short burst”, the map with the highest number of districts won (for the party in question) is then the seed for the next short burst of size 10. (If there is more than one map with the same highest number of districts won among those 10, the last such produced map is the seed for the next short burst.) This process is repeated thousands of times in an effort to find the maximum number of majority-minority districts achievable. Note that the Short Burst method can be thought of as a non-deterministic version of beam search on the metagraph [^15] The “metagraph” here is the graph whose nodes are maps. Two maps are adjacent if one can be obtained by the other through a single step in the Markov Chain constructing the ensemble of maps. In this case, that is through a single ReCom move in GerryChain. of redistricting maps, using depth-first search instead of breadth-first, and in which a random selection of maps near to the current map are explored. This process is repeated for each state/map under consideration, and each party. We considered 6 states, 3 maps per state (State House, State Senate, Congressional districting maps), and 2 parties (Democratic and Republican). This made for a total of 36 runs. We first describe the states chosen and then describe the parameters for each Short Burst run.
Our choice of which states to evaluate was based on two main factors. Firstly, we needed to use states for which there was reliable data on which we could run GerryChain’s ReCom MCMC process. The Metric Geometry Gerrymandering group [31] has an excellent GitHub page with this data for some states, and we chose our states from among those that they had available, which all used maps from the 2011 redistricting cycle. Our studies are not intended to focus on any particular map, or on the specific election for that state (which is used as a proxy for party preference), but rather to focus on the metrics themselves. Thus, we are unconcerned that these states’ maps and partisan data are 6-10 years old.
Secondly, we wanted to choose states with different types of political geographies on which to test the Mean-Median Difference and Partisan Bias. Some states are fairly homogeneous, with most parts of the state having the same partisan breakdown. Other states have pockets that more heavily lean Republican and other pockets that more heavily lean Democratic; such states may have an overall clear partisan lean (Democratic or Republican) or they may be more “purple.” The states that we chose are Massachusetts, Michigan, Oklahoma, Oregon, Pennsylvania, and Texas, which we believe represents a variety of political geographies [^16] These are the same states chosen for a similar analysis, in [45].. We include in Figure 9 a choropleth of each state (using the election data that we used for our analysis) for the reader’s reference.
Figure 9: Choropleths of the states in our analysis, using the election results from Table 7. They are, in clockwise order from the top left, Massachusetts, Oklahoma, Pennsylvania, Oregon, Texas, and Michigan.
State MA OK PA MI TX OR
Election Used 2018
US Senate
2018
Governor
2016
US Senate
2016
Presidential
2014
US Senate
2018
Governor
Table 7: Elections chosen as a proxy for partisan support for each state.
For congressional maps, we allowed up to 5% deviation from ideal population in each district. For the lower and upper house maps, we allowed up to 11% deviation from ideal population. As noted in [18], “Excessively tight requirements for population balance can spike the rejection rate of the Markov chain and impede its efficiency or even disconnect the search space entirely.” In [18], they use a deviation of 5% for their congressional map of Arkansas, which is why we chose 5% for our congresssional maps here. State legislatures have significantly looser restrictions on population deviation; Brown v Thomson [6] states that state legislative districts can deviate by up to 10% in population, and many states do have population deviations of that magnitude. We allowed a deviation of up to 11% because this is sufficiently close to 10% and is needed in order for the Markov Chain process to not become quickly stuck and be unable to produce additional sample maps.
As described in [18], the ReCom process in GerryChain naturally samples plans with more compact districts with higher probability. Outside of this built-in preference for compact maps, we did not impose additional compactness constraints.
For each state, each party, and each districting map, we ran 10 short burst samples, each of size 5,000, and each having burst length of size 10. The burst length of size 10 was theorized to be the ideal burst length in [10], which also found that the exact value of the burst length did not drastically affect the results. The step size of 5,000 was chosen due to the fact that this length corresponded to a drastic slowdown in the increase of the metric value in the study from section 5.2.3 of [10], and thus reaches close to the maximum possible number of districts won. We ran these bursts of 5,000 steps 10 times as was done by Cannon et al. in their study showing the efficacy of the Short Burst method [10], in order to validate the consistency of the results. This entire process produced potentially 45,000 different maps [^17] Since the last maximal map of each burst is the first map in the next burst, but the seed at the beginning of each sample is the same. for all of the state, party, and map combinations. These maps provide a set of examples of ‘locally optimized’ plans that can serve as a proxy for maps that are expected to return a large number of seats for a particular party, using the actual political geography of each state.
We note that the authors of [45] also used this same kind of Short Burst analysis to test the “gameability” of some metrics intended to detect gerrymandering (specifically, the GEO Metric, Mean-Median Difference, Declination, Efficiency Gap). They define a metric as “gameable” if the metric value can stay within acceptable bounds [^18] For a metric $m$, they considered “acceptable bounds” to be $0.16\inf(m)\leq m\leq 0.16\sup(m)$, based on the bounds recommended by the creators of the Efficiency Gap[46], and the apples-to-apples conversion of those bounds to the other metrics., while allowing a map with the most extreme number of Democratic-won (or Republican-won) districts that can be achieved by that map. The authors of [45] allowed us to use the same code to produce the Figures in Section 5.2
5.2 Results of Empirical Study
As mentioned in Section 5.1, we have results for both the Democratic party and the Republican Party for three different maps on each of six states. Note that, as explained in Section 5.1, the method we use to produce our maps is a hill-climbing method recently developed by Cannon et al [10] which utilizes the GerryChain library. Thus, the maps under consideration are not constructed in the same way as those constructed for a typical ensemble analysis (see Section 5.1 for details) and are not intended to provide representative samples from a specified distribution. Instead, they represent a large set of examples that have been optimized for extreme seats outcomes.
Recall that, if a metric is to be able to detect maps with an extreme number of districts won, then it should have more extreme values on extreme maps. In other words, if we see the metric values for maps with $k$ districts won by a particular party, that range of metric values should generally go up as $k$ goes up [^19] Or the metric value should go down as $k$ goes down; the direction depends on whether $V$ and $S$ are measured for the Democratic party of Republican party. For our calculations, we always let $V$ represent the Democratic party’s vote share and $S$ the Democratic party’s seat share. Thus, for the Democratic party we would expect the metric value to go up as $k$ (the number of seats won by Democrats) goes up, while for the Republican party we would expect the metric value to go down as $k$ (the number of seats won by Republicans) goes down..
We see sample results in Figures 10 and 11 for Massachusetts’ State House. Just to compare the Mean-Median Difference and Partisan Bias to other gerrymandering metrics, we include the Efficiency Gap [46] and the GEO Metric ratio[8] [^20] Since these metrics are used purely for comparison purposes, we do not define them here. Their definitions are not our concern; our concern is simply that other metrics intended to detect gerrymandering do have extreme values on more extreme maps, as one would expect. The Efficiency Gap is defined in [46], the GEO metric is defined in [8], and the GEO metric ratio is defined in [45] .